OpenStudy (anonymous):

7.2-3). Multiply the following expression. (x^2+3x-10) x+1/x-2

OpenStudy (watchmath):

\((x+5)(x-2)\cdot\frac{x+1}{x-2}=(x+5)(x+1)\)

OpenStudy (anonymous):

hey